### v1

# DISC: Dynamic Decomposition Improves LLM Inference Scaling

Abstract
Inference scaling methods often rely on decomposing problems into steps (or groups of tokens), followed by sampling and selecting the best next steps. However, these steps and their sizes are often predetermined or manually designed based on domain knowledge. We propose dynamic decomposition, a method that adaptively and automatically fractions solution and reasoning traces into manageable steps during inference. By more effectively allocating compute – particularly through subdividing challenging steps and prioritizing their sampling – dynamic decomposition significantly improves inference efficiency. Experiments on benchmarks such as APPS, MATH, and LiveCodeBench demonstrate that dynamic decomposition outperforms static approaches, including token-level, sentence-level, and single-step decompositions. These findings highlight the potential of dynamic decomposition to improve a wide range of inference scaling techniques.
Machine Learning, LLMs, Inference scaling, test-time compute, planning, reasoning, decomposition
1 Introduction
Scaling inference efficiency remains a fundamental challenge for large language models (LLMs). Many existing approaches improve inference by decomposing problems into smaller steps and systematically exploring different solutions (Feng et al., 2023; Zeng et al., 2024; Wu et al., 2024; Nori et al., 2024; Snell et al., 2024; Brown et al., 2024; Gandhi et al., 2024; Lee et al., 2025; Light et al., 2024a; Wang et al., 2025).
Some decomposition methods rely on domain-specific heuristics and hand-crafted rules (Yao et al., 2024; Zelikman et al., 2023; Zhou et al., 2022). However, manually partitioning problems or designing task-specific heuristics is costly and lacks generalization. Moreover, identifying critical steps for an LLM can be non-trivial for humans. As shown in Sec. 3.5, LLMs may assign importance to seemingly trivial words (e.g., therefore or which), which, while counterintuitive to humans, play a crucial role in autoregressive generation (Lin et al., 2025). Other approaches employ fixed, uniform step sizes, such as token- or sentence-level decomposition (Feng et al., 2023; Guo et al., 2025). All these methods rely on static decomposition strategies, where step sizes are predefined or determined via heuristics. Such rigidity wastes compute on steps that are easy for the LLM (but potentially difficult for humans) while undersampling more challenging steps.
To overcome these limitations, we propose DISC (Dynamic decomposition Improves Scaling Compute), a recursive inference algorithm that dynamically partitions solution steps based on difficulty. Unlike prior methods, DISC adapts decomposition granularity during inference based on both the available budget and problem complexity, ensuring finer granularity for more difficult steps. By leveraging the autoregressive nature of LLMs, DISC efficiently locates difficult steps through binary partitioning, focusing compute on challenging regions rather than wasting resources on trivial steps. DISC is generalizable and requires no human supervision, domain-specific heuristics, prompt engineering, or process annotations, making it widely applicable across tasks.
Our main contributions are:
  • We introduce DISC, a method for recursive partitioning and decomposing solutions during inference without human supervision, domain-specific heuristics, or process reward models.
  • We demonstrate how DISC integrates decomposition with inference-time search, allocating compute to high-impact, difficult steps.
  • We show that DISC improves inference scaling in terms of both sample efficiency and token efficiency.
  • We provide insights into how LLMs reason and plan by identifying critical steps in their generation process.
2 Preliminaries
Figure 1: Comparison of different automatic decomposition methods based on step size determination.
2.1 Problem Setting
We consider a reasoning and code generation setting where a dataset $\mathcal{X}=\{\boldsymbol{x}^{(i)}\}_{i=1}^{N}$ consists of problem prompts $\boldsymbol{x}$, and a reward model $R:\mathcal{X}\cdot\mathcal{Y}\rightarrow[0,1]$ evaluates generated solutions $\boldsymbol{y}\in\mathcal{Y}$. This includes program synthesis, where correctness is verified using ground-truth tests (Chen et al., 2021; Austin et al., 2021), and mathematical reasoning, where solutions are validated numerically (Hendrycks et al., 2021a; Cobbe et al., 2021). The reward model can be a ground-truth verifier, a trained heuristic (Zhang et al., 2024), self-consistency (Wang et al., 2023a), or an LLM-as-a-judge (Zheng et al., 2023). Since our focus is on step decomposition rather than verification, we use the ground-truth reward model where available. We assume access to a pretrained language model $\pi$ that generates text autoregressively. A generated response $\boldsymbol{y}$ consists of both the final solution and the reasoning chain leading to it, and can be represented as a sequence of tokens $\boldsymbol{y}=(y_{0},...,y_{L_{\boldsymbol{y}}})$. Additionally, solutions can be partitioned into solution steps $\boldsymbol{y}=(\boldsymbol{y}_{0},...,\boldsymbol{y}_{K})$, where each step $\boldsymbol{y}_{i}$ is a contiguous string of tokens. A partial solution up to step $k$ is defined as $\boldsymbol{y}_{1...k}:=\boldsymbol{y}_{1}\cdot\boldsymbol{y}_{2}\cdot...\cdot\boldsymbol{y}_{k}$, and its rollout or completion, denoted $\boldsymbol{y}_{1...k+}$, is the continuation generated by $\pi$ until an end-of-sequence token (EOS). The size of a solution step, $|\boldsymbol{y}_{i}|$, refers to its length in tokens or characters.
2.2 Prior Automatic Decomposition Methods
Single-step generation. In a single-step generation, the entire solution is generated in one pass from the prompt to the EOS token, treating it as a single action. This approach underlies the widely used inference scaling method best of n (BoN) (Cobbe et al., 2021; Lightman et al., 2023; Snell et al., 2024; Liang et al., 2024), where $n$ complete solutions are sampled, and the highest-scoring one is selected. Single-step generation also plays a role in alignment and fine-tuning methods such as DPO (Rafailov et al., 2024) and RLOO (Ahmadian et al., 2024).
Token-level decomposition. At the opposite end of the spectrum, token-level decomposition treats each atomic token as an individual step. While this approach dramatically increases search complexity, it enables fine-grained search that can yield higher performance gains given sufficient compute (Feng et al., 2023).
Newline and sentence-level decomposition. A commonly used decomposition method segments LLM generations into sentences or lines based on delimiters such as periods or newlines (Hao et al., 2023; Feng et al., 2023; Yao et al., 2024). Typically, each newline corresponds to a new paragraph, equation, or line of code, which often encapsulates a distinct reasoning step.
Problem: Automatic and scalable decomposition Existing decomposition methods are task-specific, manual, and static, limiting their adaptability and scalability.
2.3 Step Sampling
Inference-time scaling methods must balance exploration at the current step with exploration of future steps. We implement a simple dynamic sampling process, referred to as negative binomial sampling, where we continue sampling completions until the sum of their rewards exceeds a predefined threshold $\sigma$. More formally, the number of samples $M$ drawn from a partial solution $\boldsymbol{y}_{1...k}$ is the smallest integer satisfying $\sum_{i=1}^{M}R(\boldsymbol{x}\cdot\boldsymbol{y}_{1...k+}^{(i)})\geq\sigma$.
Where $\boldsymbol{y}_{1...k+}^{(i)}$ represents the $i$-th sampled completion from the partial solution. This process ensures efficient allocation of compute by dynamically adjusting the number of samples per step. It achieves this by either: (a) continuing to sample until a sufficiently high-reward completion is found, or (b) stopping early when additional sampling is unlikely to yield significant improvements, thus redirecting the compute to future steps. The stopping criterion is governed by $\sigma$: when accumulated reward from completions surpasses $\sigma$, the method assumes further sampling is unnecessary. For a fair comparison, we apply this sampling method uniformly across all decomposition methods in our experiments.
2.4 Inference Scaling Methods and Decomposition
Since our study focuses on decomposition rather than search, we primarily use greedy step search as the search method. In greedy step search, multiple candidate steps are sampled at each iteration, but only the highest-scoring step is retained, while the rest are discarded. The process then repeats, conditioning future steps on the best step found so far. We also perform ablation studies comparing Monte Carlo Tree Search (MCTS)  (Feng et al., 2023; Light et al., 2024b) and beam search  (Xie et al., 2024), two commonly used inference scaling methods. These comparisons, presented in Sec. 4.2, highlight how different search strategies interact with decomposition. Additional details on MCTS and beam search are provided in App. E.
3 Methodology
Figure 2: DISC recursively partitions the most challenging step – i.e., the one with the lowest priority metric $h$ – to progressively refine and find the optimal solution.
3.1 DISC Algorithm
The DISC algorithm employs recursive binary decomposition to iteratively break down complex solutions into smaller, more manageable steps. Given a problem prompt $\boldsymbol{x}$, the algorithm outputs a decomposition of a solution, $\boldsymbol{y}=(\boldsymbol{y}_{1},\boldsymbol{y}_{2},\dots,\boldsymbol{y}_{K})$, such that the concatenation $\boldsymbol{y}_{1...K}$ forms a complete solution to $\boldsymbol{x}$.
The algorithm operates in two key stages:
1. Solution sampling. Starting from $\boldsymbol{x}$ and $\boldsymbol{y}_{0}=\emptyset$, the algorithm generates complete solutions $\boldsymbol{y}_{k}\sim\pi(\cdot|\boldsymbol{x}\cdot\boldsymbol{y}_{1...(1-k)})$ using a policy $\pi$ like in single step generation. The best solution, $\boldsymbol{y}^{*}_{k}$, is selected based on the reward model $R$.
2. Recursive partitioning. The selected solution $\boldsymbol{y}^{*}_{k}$ is partitioned into two segments, $\boldsymbol{y}^{*}_{k}=\boldsymbol{y}^{*}_{a}\cdot\boldsymbol{y}^{*}_{b}$, based on a predefined partition fraction $\alpha$, where $|\boldsymbol{y}^{*}_{a}|\approx\alpha|\boldsymbol{y}^{*}_{k}|$. For each part, a priority metric $h$ is estimated: $\hat{h}(\boldsymbol{y}^{*}_{a}|\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})$ and $\hat{h}(\boldsymbol{y}^{*}_{b}|\boldsymbol{x}\cdot\boldsymbol{y}^{*}_{1...a})$, usually through rollouts of the step using $\pi$. The part with the lower priority is further partitioned.
  • If $\hat{h}(\boldsymbol{y}^{*}_{a}|\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})\geq\hat{h}(\boldsymbol{y}^{*}_{b}|\boldsymbol{x}\cdot\boldsymbol{y}^{*}_{1...a})$, additional samples are sampled for $\boldsymbol{y}^{*}_{b}$, with the process repeating on the new best solution, $\boldsymbol{y}^{\prime*}_{b}$.
  • Conversely, if $\hat{h}(\boldsymbol{y}^{*}_{a}|\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})<\hat{h}(\boldsymbol{y}^{*}_{b}|\boldsymbol{x}\cdot\boldsymbol{y}^{*}_{1...a})$, the first segment $\boldsymbol{y}^{*}_{a}$ is further partitioned. The first step corresponds to the $\alpha$ fraction of $\boldsymbol{y}^{\prime*}_{a}$, with the remaining part of the full solution forming the second step.
This recursive process is illustrated in Fig. 2. The pseudocode for DISC is provided in Alg. 1, with an annotated Python implementation in App. A.
The priority metric $h$ serves as the central heuristic for determining which solution steps to prioritize. It estimates the “difficulty” or “potential for improvement” of a step $\boldsymbol{y}_{k}$ given the context $\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)}$, computed via rollouts of the policy $\pi$. Specifically, $h(\boldsymbol{y}_{k}|\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})$ is estimated by sampling continuations and evaluating their outcomes.
In practice, estimating $h$ and generating new samples occur simultaneously, as both rely on rollout-based computations (Sec. 3.2). Unlike standard decomposition methods, DISC does not process steps in strict temporal order, resembling goal-directed planning (Parascandolo et al., 2020) and backtracking.
Intuition and Benefits. DISC partitions difficult steps into smaller, simpler sub-steps, allocating additional resources to refine them. This approach is particularly effective for inference scaling, where computing must be used judiciously. A binary decomposition strategy enables fast identification of difficult or high-potential steps.
Key Insight: Recursive partitioning Top-down, recursive partitioning means that we can both efficiently locate critical steps and also dynamically determine step sizes based on our budget.
Dynamic Compute Allocation. A key advantage of DISC is its ability to prioritize challenging or high-potential steps, improving solution quality while minimizing compute waste. By iteratively refining steps with low priority scores, DISC adaptively allocates more resources to difficult steps and less to simpler ones, optimizing inference efficiency.
Key Insight: Adaptive compute allocation DISC dynamically allocates inference compute to harder steps, optimizing solution quality and resource efficiency.
3.2 Priority Metric
We consider two intuitive priority metrics for step selection: Q-value priority and Z-score priority, which are visualized in Fig. 4.
Q-value based priority (DISC-Q). Given a partial solution $\boldsymbol{y}^{*}_{1...k}=\boldsymbol{y}^{*}_{1...(k-1)}\cdot\boldsymbol{y}^{*}_{a}\cdot\boldsymbol{y}^{*}_{b}$, we aim to prioritize either $\boldsymbol{y}^{*}_{a}$ or $\boldsymbol{y}^{*}_{b}$ for refinement. The core intuition behind Q-value prioritization is that steps with lower Q-values indicate areas needing refinement, directing compute toward the most challenging parts of the solution. More formally, we define the Q-priority metric under policy $\pi$ as:
$h_{Q}(\boldsymbol{y}^{*}_{k}\mid\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})=\mathbb{E}_{\boldsymbol{y}_{k}}\left[Q^{\pi}(\boldsymbol{y}_{k}\mid\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})\right].$
Here, $\boldsymbol{y}_{k}$ represents alternative steps sampled from $\pi$, and $Q^{\pi}(\boldsymbol{y}_{k}\mid\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})$ denotes the Q-value of $\boldsymbol{y}_{k}$, conditioned on the partial solution $\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)}$. Equivalently, $h_{Q}(\boldsymbol{y}^{*}_{k})$ can be interpreted as the value function $V^{\pi}(\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})$, where $V^{\pi}$ represents the expected reward achievable from the given partial solution. The expectation is taken over sampled candidates $\boldsymbol{y}_{k}\sim\pi(\cdot\mid\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})$, constrained by $|\boldsymbol{y}_{k}|=|\boldsymbol{y}^{*}_{k}|$. This metric helps identify difficult steps that the LLM is likely to get wrong, guiding partitioning toward the most critical refinements.
To estimate $h_{Q}$ for the second step $\boldsymbol{y}^{*}_{b}$, we sample $\boldsymbol{y}_{b}$ from $\pi$, generate rollouts $\boldsymbol{y}_{(k+1)+}$, compute rewards, and average the outcomes. For the previous step $\boldsymbol{y}^{*}_{a}$, we reuse rollouts from earlier partitioning, as the mean of these previously generated samples provides an unbiased estimate of the Q-priority metric:
$h_{Q}(\boldsymbol{y}^{*}_{a}\mid\boldsymbol{x}\cdot\boldsymbol{y}^{*}_{1...(k-1)})=\mathbb{E}\left[R(\boldsymbol{y}^{*}_{1...(k-1)+})\right].$
By leveraging existing rollouts, DISC avoids redundant sampling, improving computational efficiency. Once the rollouts for $\boldsymbol{y}^{*}_{(k+1)}$ are available, the best completion $\boldsymbol{y}^{*}_{(k+1)+}$ is selected as the next step to partition. This dual use of rollouts optimizes both metric estimation and inference sampling.
Combining priority estimation and test-time search By integrating LLM policy sampling for both metric estimation and search, we can significantly enhance computational efficiency.
Figure 3: Priority metric estimation. We can estimate the priority metric $\hat{h}$ using Monte Carlo rollouts of the LLM policy $\pi$ for each step.
Z-score based priority (DISC-Z). To allocate more compute to steps with higher potential for improvement, we estimate the probability of sampling a better step given existing samples. Assuming the Q-values of sampled steps follow a normal distribution, we model this probability using the cumulative distribution function (CDF). Given a mean $\mu_{r}$ and standard deviation $\sigma_{r}$ of sampled Q-values, the probability of sampling better than the best-observed step $\boldsymbol{y}^{*}_{k}$ with Q-value $q^{*}$ is:
$1-\operatorname{CDF}\left(\frac{q^{*}-\mu_{r}}{\sigma_{r}}\right)=1-\operatorname{CDF}(z^{*}),$
where $z^{*}$ is the Z-score. Since the CDF is monotonic, we compare steps based on their Z-scores.
We formally define the Z-score priority metric as:
$h_{Z}(\boldsymbol{y}^{*}_{k}\mid\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})=\frac{q^{*}-\mathbb{E}[Q^{\pi}(\boldsymbol{y}_{k}\mid\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})]}{\operatorname{Std}[Q^{\pi}(\boldsymbol{y}_{k}\mid\boldsymbol{x}\cdot\boldsymbol{y}_{1...(k-1)})]}.$
Since $q^{*}-\mu_{r}$ represents the advantage of step $\boldsymbol{y}^{*}_{k}$, the Z-score metric can be interpreted as a standard deviation-scaled advantage. Lower $h_{Z}$ values indicate steps with greater room for improvement, guiding decomposition toward those with higher variance in performance.
Figure 4: Different priority metrics. Sampled rewards of different steps can be visualized as a distribution.
3.3 DISC and Search Methods
DISC can also be used to enhance Monte Carlo Tree Search (MCTS) and other inference scaling and search methods. Recall that in our partition step, we greedily partition the best solution step $\boldsymbol{y}^{*}$. Instead of greedily partitioning the best step, we can partition the top $k$ best steps instead, and select which step to partition using the upper confidence tree (UCT) formula. We can also use beam search to prune out steps we do not want to partition further. We explain MCTS and beam search in detail in App. E and present results of combining DISC with search in Sec. 4.9.
Takeaway: Dynamic step sizes can improve search DISC can enhance search based inference scaling methods by determining what step size to search across.
3.4 A Motivating Example on DISC-Z
We use the Wiener process $W(t)$ as an example where there are intractably many actions and steps. Suppose we start at $t=0$ with $W(0)=0$. At each round $k$, the algorithm can choose one of the two options:
  1. 1.
    samples a trajectory and observe the final value $W(T)$ at time $t=T$, as the reward signal. Denote the whole trajectory as $w_{k}(\cdot)$.
  2. 2.
    chooses one trajectory from the previous rounds (denoted as $w_{s}(t)$ for round $s$), and time $t_{0}$; then sample a trajectory at $t=t_{0}$ with $W(t_{0})=w_{s}(t_{0})$. Denote the concatenated trajectory as $w_{k}(\cdot)$ with $w_{k}(t)=w_{s}(t)$ when $t\leq t_{0}$.
Note that we are only able to observe the final reward $W(t)$. At any intermediate time $t\in(0,T)$, the current value $W(t)$ is not observable. The goal is to design an algorithm that can reach the highest reward among the $K$ trajectories. Formally speaking, we aim to maximize the maximum:
$\max_{k\in K}w_{k}(T).$
One naive solution is to call option 1 for $K$ times and return the best-of-$K$ reward, each following:
$W(T)$ $\sim\mathcal{N}(0,T).$
Alternatively, suppose there is a promising path $w(\cdot)$ with a high final reward $w(T)=R$. It is natural to consider starting at some midpoint $\alpha T$ ($10<\alpha<$ 1) and perform more completions to obtain an even higher reward than $R$. The reward distribution sampled this way is
$W^{\prime}(T)\sim\mathcal{N}(w(\alpha T),(1-\alpha)T).$
The remaining question is which $\alpha$ we should choose. One option is to maximize the probability that the newly sampled reward is higher than $R$:
$\mathbb{P}(W^{\prime}(T)>R)=1-\Phi\bigg{(}\frac{R-w(\alpha T)}{\sqrt{(1-\alpha)T}}\bigg{)}.$
3.5 Example Decomposition
With sample budget 100, the decomposition of a MATH problem is as follows, where color indicates the value of the priority metric $h$ of each step (yellow low, purple high).
DISC example decomposition Let the length of the rectangle be $l$ and the width of the rectangle be $w$. Since the perimeter of the rectangle is 24 inches, we have that $242l+2w=$ 24, so $l+w=12$. We wish to maximize the area of the rectangle, which is $A=lw$. Let $l=12-w$ and plug into the area: ${\color[rgb]{0.265625,0.00390625,0.328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.00390625,0.328125}A=(12-w)w\Rightarrow\qquad A=12w-w^{2}}$ Now, we differentiate $A$ with respect to $w$: ${\color[rgb]{0.265625,0.00390625,0.328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.00390625,0.328125}A^{\prime}(w)=12-2w}$ We wish to maximize $A$, so we set $A^{\prime}(w)=0$, and solve for $w$: ${\color[rgb]{0.265625,0.00390625,0.328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.00390625,0.328125}12-2w=0\Rightarrow\qquad w=6}$ Since $l=12-w$, we have that $l=12-6=6$. Therefore, the area of the rectangle is $A=lw=6\cdot 6=\boxed{36}$.
Once the LLM generates the first three steps, the rest is easy. Interestingly, ‘which’ is an important decision point which helps decide how the LLM will complete the solution.
Takeaway: Autoregressive models require autoregressive decomposition While words such as ‘which’, ‘therefore’, etc. may not seem like important steps to humans, they actually represent important steps for autoregressive LLMs which are trained on next token prediction.
4 Experimental Results
4.1 Benchmarks
We evaluate DISC on three benchmarks: APPS, MATH, and LiveCodeBench, to assess its impact on inference scaling for both coding and reasoning.
APPS (Hendrycks et al., 2021a) consists of 5000 competitive programming problems across three difficulty levels, with the competition-level subset being the hardest. We evaluate on a 200-problem subset due to computational constraints. MATH (Hendrycks et al., 2021b) comprises 12,500 math problems. Since the ground-truth verifier provides only binary rewards, we use a pretrained ORM (Xiong et al., 2024), trained via the method in (Wang et al., 2024b), with Llama-3.1-8B-Instruct as the base model. We test on a 500-problem subset (MATH500), identical to prior work (Wang et al., 2024b; Lightman et al., 2023). LiveCodeBench (Jain et al., 2024) is a continuously updated dataset from Leetcode, AtCoder, and CodeForces, ensuring LLMs have not been exposed to test problems. We evaluate on the 108 problems uploaded between 10/01/2024 and 12/01/2024 to prevent contamination.
Figure 5: Token-level comparison on APPS competition level. DISC consistently outperforms prior decomposition methods in efficiency.
Figure 6: Token-level comparison on MATH500. DISC outperforms other decomposition methods, demonstrating its effectiveness in mathematical reasoning.
Figure 7: Token-level comparison on LiveCodeBench. DISC achieves superior inference scaling over TokenSplit, LineSplit, and BoN.
4.2 Decomposition Comparison
We compare DISC against three prior decomposition methods from Sec. 2.2: TokenSplit (token-level decomposition), LineSplit (newline-based decomposition), and BoN (treating the entire solution as a single step).
Across all benchmarks, DISC achieves superior scaling and performance under both fixed token budgets (Fig. 7) and sample budgets (Fig. 12). We evaluate two key metrics: Pass@k, the proportion of problems solved within a sample budget $k$, and Pass@token, the proportion solved within a given token budget. Notably, DISC consistently outperforms static decomposition methods on APPS, MATH, and LiveCodeBench (Fig. 567), demonstrating its ability to allocate compute adaptively for improved inference efficiency. Extended results and analyses for each benchmark are provided in App. D.1D.4, and D.5.
4.3 Decomposition Analysis and Interpretation
Our results strongly indicate that decomposition—whether line-based, token-based, or DISC —improves sample quality. Figures 8 and 9 illustrate how the mean and variance of sampled rewards evolve with the step number, which represents the order in which a step is explored. Higher step numbers correspond to deeper search levels, where solutions are partitioned into finer-grained steps. As shown in Fig. 8, increasing step number correlates with higher-quality solutions, demonstrating that finer-grained decomposition improves sample quality. Additionally, Fig. 9 shows that reward variance decreases as step count increases, highlighting how decomposition enhances sampling precision.
Furthermore, DISC achieves better performance with fewer partitions under a fixed sampling budget (Fig. 10). We distinguish between actual partitions, the number of steps effectively explored, and planned partitions, the number of partitions intended by the method. Token and line split methods generate a large number of planned partitions (Fig. 10) but search over at most 15 steps due to budget constraints. In contrast, DISC dynamically adjusts the number of partitions based on available budget, efficiently identifying and focusing on critical steps.
Takeaway: Decomposition and sample quality Finer-grained decomposition improves sampled solution quality and reduces reward variance.
Figure 8: Average reward per step, averaged over APPS problems. From step 3 onward, higher step counts strongly correlate with increased average reward, demonstrating the effectiveness of decomposition. The dip between steps 1 and 3 likely occurs because simple problems are solved early, preventing further search.
Figure 9: Standard deviation of rewards per step, averaged over APPS problems. Decomposition reduces sampling variance, improving precision at deeper search depths.
Figure 10: Comparison of actual and planned partitions on APPS. DISC outperforms other methods with fewer partitions by efficiently identifying critical steps. Unlike token and line split methods, which plan many partitions but search only a subset, DISC dynamically adjusts partitioning based on budget.
4.4 Interaction Between Temperature and DISC
Figure 11: Effect of temperature on DISC performance on APPS with gpt-4o-mini. Unlike BoN and other inference scaling methods, DISC achieves higher performance at lower temperatures.
We perform ablation studies to analyze the impact of temperature on DISC. Typically, inference scaling methods achieve optimal performance at temperatures around 0.6–0.8, as increased temperature promotes sample diversity (Wang et al., 2024a). Surprisingly, however, DISC performs better at lower temperatures, as shown in Fig. 11. This trend is in stark contrast to BoN (Fig. 22), where higher temperatures are generally beneficial.
We believe this phenomenon arises because DISC depends on accurately estimating the priority metric $h$ at each step. Lower temperatures reduce sample variance, leading to more reliable estimates of $h$, which in turn improves step selection. This is further supported by Fig. 20, which shows that lower temperatures yield lower standard deviations per step, indicating increased sampling consistency. Additional details and analyses can be found in App. C.1.
4.5 Self-Generated Validation Tests
We also evaluate DISC in a more practical setting where a ground-truth reward model is unavailable for code generation  (Chen et al., 2022, 2023b; Zhou et al., 2024). Instead of relying on predefined test cases, we prompt the LLM to generate validation test cases based on the problem prompt. In real-world applications, manually curated ground-truth test cases are often costly to obtain, making self-generated validation a more scalable approach. The results, shown in Fig. 12, indicate that DISC continues to scale better than other methods in this setting. Additional results and details are provided in App. D.3.
Figure 12: Comparison of Pass@k performance on APPS with ground truth tests (right) and self generated validation tests (left) using gpt-4o-mini. DISC scales more effectively in both.
4.6 Ablation on Base LLM Model
We evaluate DISC across different LLMs, including open-source models. As shown in Fig. 13 and Fig. 27, DISC significantly enhances performance even for weaker models. Specifically, it improves Llama’s pass rate from 1% to 5.5%, a 550% relative increase, and Mistral’s from 0% to 3.5%, demonstrating substantial gains even from a nonzero baseline. Additional details and analyses are in App. C.3.
Figure 13: Pass@k scaling curve for open-source models with DISC on APPS. DISC substantially improves performance across different LLMs, including Llama and Mistral.
4.7 Interaction Between Priority Metric and DISC
We conduct an ablation study to examine how the choice of priority metric affects DISC performance. In addition to the Q-based and Z-based priority metrics (DISC-Q and DISC-Z) introduced in Sec. 3.2, we evaluate three baselines: DISC-R (random step selection), DISC-negQ, and DISC-negZ (which prioritize the opposite steps of DISC-Q and DISC-Z, respectively). As shown in Fig. 14, the selection of a priority metric significantly impacts performance. Both DISC-Q and DISC-Z significantly outperform random selection and their inverse counterparts, demonstrating the effectiveness of their priority heuristics. Additional details and analysis are in App. C.2.
Figure 14: Effect of priority metrics on DISC performance on APPS with gpt-4o-mini. Both Q-based and Z-based priority metrics outperform random selection and their inverses, highlighting their effectiveness.
4.8 Ablation on Partition Fraction $\alpha$
We conduct an ablation study to analyze the effect of the partition fraction $\alpha$ on DISC performance. As shown in Fig.  15 and  31, the optimal range appears to be $0.15\leq\alpha\leq 0.25$. Lower partition fractions ($\alpha<0.5$) tend to perform better due to the asymmetric cost of sampling from different halves of the partition. Sampling from the first half requires generating more tokens, while the second half requires fewer, making it crucial to partition the first half more conservatively. Additional analysis are in App. C.4.
Figure 15: Pass@k comparison of different DISC partition fractions $\alpha$ on APPS competition level. The range $0.15\leq\alpha\leq 0.25$ appears optimal.
4.9 Search and DISC
We demonstrate that search methods such as MCTS and beam search can be combined with DISC. As shown in Fig. 47 in the Appendix, greedy search explores deeper partitions given the same search budget due to its greedy nature, while MCTS and beam search reach similar, shallower depths. However, MCTS allocates the search budget more effectively than beam search, leading to higher performance, as seen in Fig. 16. Additional details and analysis are in App. E.
Figure 16: Pass@k comparison of search methods combined with DISC on APPS with gpt-4o-mini. MCTS scales best, followed by greedy search, then beam search (with beam size 2).
5 Related Work
Inference scaling. Inference scaling has emerged as a dominant paradigm, driven by the introduction of o1- and r1-like chain-of-thought reasoning models (Snell et al., 2024; Brown et al., 2024; Manvi et al., 2024; Leea et al., 2025). Several works examine the trade-off between inference compute and training compute (Guan et al., 2025; Chen et al., 2024b). LLM inference often relies on decomposing complex problems into intermediate reasoning steps, as seen in chain-of-thought (CoT) prompting (Wei et al., 2022; Sprague et al., 2024; Wang & Zhou, 2024) and its variants (Kojima et al., 2022; Zhou et al., 2023; Wang et al., 2023b; Li et al., 2023). We extend inference scaling by introducing a new approach for adaptive compute allocation (Manvi et al., 2024).
LLM reasoning and code generation. LLM reasoning and code generation are central tasks for inference scaling. Evolutionary inference scaling methods have been explored in program generation (Liventsev et al., 2023; Chen et al., 2023a; Romera-Paredes et al., 2024; Lehman et al., 2023; Hemberg et al., 2024). Domain-specific decomposition strategies have been applied in code generation, such as function-based decomposition (Chen et al., 2024a; Zenkner et al., 2024; Levin et al., 2025). More broadly, decomposition often involves prompting LLMs to generate subtask completions (Hernández-Gutiérrez et al., 2024; Khot et al., 2022; Dua et al., 2022), which differs from methods that refine a single LLM generation.
Reinforcement learning and Monte Carlo methods. Unlike standard RL, our setting resembles a search problem where the goal is to identify the single highest-reward path. Cazenave (2009) demonstrated that nested Monte Carlo search can accelerate optimal pathfinding. Under the bandit setting, this can be formulated as identifying the arm with the highest maximum reward rather than the highest mean reward (Cicirello & Smith, 2005; Carpentier & Valko, 2014).
6 Conclusion
We introduce DISC, a dynamic decomposition framework that adaptively partitions solution steps based on difficulty, improving inference scaling by directing compute toward critical steps while balancing exploration and resource allocation. DISC seamlessly integrates with search-based methods such as MCTS and beam search, further enhancing performance. It also identifies challenging steps for LLMs, aiding curriculum learning, fine-tuning, and dataset augmentation. By dynamically adjusting partitioning based on available compute, DISC enables more adaptive and efficient reasoning in large language models, with broad implications for both training and inference optimization.
7 Impact Statements
This work on dynamic decomposition aims to advance the field of LLM by improving the efficiency and accuracy of large-scale inference methods, particularly in domains like code generation, mathematical problem-solving, and automated reasoning. By dynamically allocating computational resources to challenging reasoning steps, this approach has the potential to enhance AI-powered tools in software development, STEM education, and scientific research. While the core contributions of this paper focus on technical advancements, we acknowledge the broader societal implications of scaling inference techniques, including risks of over-reliance on AI in decision-making and the potential propagation of errors or biases. To mitigate these concerns, future work should emphasize robust validation, fairness, and human-in-the-loop deployment strategies, particularly in high-stakes applications. Nonetheless, no immediate or specific ethical concerns are identified in the context of this research.
Appendix A Code implementation of DISC
Python implementation of DISC def dynamic_decomposition(problem, model, reward_model, split_str, complete_solution, fraction, solution_budget, split_metric, stop_threshold=-float("inf"), stop_sum_score=1.0, stop_if_solved=False, ): """ ␣␣␣␣Decomposesthesolutionusingadynamicbinarysearchapproach ␣␣␣␣Args: ␣␣␣␣␣␣␣␣problem(Problem):Theproblemtosolve ␣␣␣␣␣␣␣␣model(Model):Themodeltouseforgeneration ␣␣␣␣␣␣␣␣reward_model(function):Therewardmodeltouseforscoring ␣␣␣␣␣␣␣␣split_str(function):Thefunctiontouseforsplittingastring ␣␣␣␣␣␣␣␣complete_solution(function):Thefunctiontouseforcompletingasolution ␣␣␣␣␣␣␣␣fraction(float):Thefractiontosplitthestring ␣␣␣␣␣␣␣␣solution_budget(int):Themaximumnumberofsolutionstogenerate ␣␣␣␣␣␣␣␣split_metric(function):Themetrictouseforsplitting ␣␣␣␣␣␣␣␣stop_threshold(float):Thethresholdtostopsplitting ␣␣␣␣␣␣␣␣stop_sum_score(float):Thesumscoretostopgeneratingcompletions ␣␣␣␣␣␣␣␣stop_if_solved(bool):Whethertostopiftheproblemissolved ␣␣␣␣""" # Initialize results and decomposition steps decomp_return = { "generated_solutions": [], "decomposition": [] } while len(decomp_return["generated_solutions"]) < solution_budget: # Combine all previous steps into an intermediate solution intermediate_solution = "".join([step["step_str"] for step in decomp_return["decomposition"]]) new_scores = [] best_solution = None best_completion = None best_score = -float("inf") sum_score = 0.0 # 1) Generate completions until we generate enough samples to estimate the split metric while sum_score < stop_sum_score: proposed_completion = complete_solution(problem, intermediate_solution, model) proposed_solution = intermediate_solution + proposed_completion decomp_return["generated_solutions"].append(proposed_solution) # Update scores proposed_score = reward_model(proposed_solution) new_scores.append(proposed_score) sum_score += proposed_score # Track the best solution if proposed_score > best_score: best_solution = proposed_solution best_score = proposed_score best_completion = proposed_completion # Stop early if problem is solved if stop_if_solved and proposed_score >= 1.0: decomp_return["decomposition"].append({"step_str": proposed_completion}) return decomp_return new_metric = split_metric(new_scores) last_metric = decomp_return["decomposition"][-1]["metric"] if decomp_return["decomposition"] else None # Determine the split target. We always split the step with the highest metric is_split_new_step = last_metric is None or new_metric >= last_metric split_target = decomp_return["decomposition"][-1]["step_str"] if not is_split_new_step else best_completion # 3) Attempt to split the target split_result = split_str(split_target, fraction) if not split_result: # If we can’t split the target, we’re done decomp_return["decomposition"].append({"step_str": best_completion, "metric": new_metric}) return decomp_return # Update decomposition based on split part1, part2 = split_result if is_split_new_step: decomp_return["decomposition"].append({"step_str": part1, "metric": new_metric}) # Stopping condition based on threshold if new_metric < stop_threshold: decomp_return["decomposition"].append({"step_str": part2}) return decomp_return else: decomp_return["decomposition"][-1] = {"step_str": part1, "metric": last_metric} return decomp_return
Appendix B Pseudocode for DISC
Algorithm 1 Dynamic Decomposition
1:  Input: Problem instance ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}x}$, reward model ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r}$, partition function ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}f}$, LLM policy model ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}\pi}$, partition fraction ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}\alpha}$, solution budget ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}B}$, priority metric ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}h}$, metric stopping precision ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}\theta}$, sampling stopping threshold ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}\sigma}$, is inference mode binference
2:  Output: Final decomposition ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}$
[…]
4:  # Decompose the solution recursively until we reach the desired precision ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}\theta}$ or run out of budget ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}B}$
5:  while $|{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}.\text{generated\_solutions}|<{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}B}$ do
6:     ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{intermediate}}}\leftarrow\text{Concatenate}([\text{step.step\_str}\ \forall\ \text{step}\in{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}.\text{decomposition}])$
[…]
8:     ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}R_{\text{new}}}\leftarrow\varnothing$vartriangleright Record rewards of completions
9:     ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}\text{best}.y_{\text{final}}}\leftarrow\text{None}$, ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}\text{best}.y_{\text{completion}}}\leftarrow\text{None}$, ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}\text{best}.r}\leftarrow-\infty$vartriangleright Track the best completion
10:     # Step 1: Generate completions until we have enough samples to estimate the splitting metric. Here we use a geometric sampling distribution
11:     while $\operatorname{sum}({\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}R_{\text{new}}})<{\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}\sigma}$ do
12:        ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{completion}}}\leftarrow{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}\pi}(\cdot|{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}x},{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{intermediate}}})$
13:        ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{proposed}}}\leftarrow{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{intermediate}}}\oplus{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{completion}}}$
[…]
15:        ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r_{\text{proposed}}}\leftarrow{\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r}({\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{proposed}}})$
16:        Append ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r_{\text{proposed}}}$ to ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}R_{\text{new}}}$
17:        if ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r_{\text{proposed}}}>{\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}\text{best}.r}$ then
18:           ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}\text{best}.y_{\text{final}}}\leftarrow{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{proposed}}}$, ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}\text{best}.y_{\text{completion}}}\leftarrow{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{completion}}}$
19:           ${\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}\text{best}.r}\leftarrow{\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r_{\text{proposed}}}$
20:        end if
21:        if $\textbf{b}_{\text{inference}}\ \textbf{and}\ {\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}r_{\text{proposed}}}=1.0$ then
[…]
23:           Return ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}$vartriangleright Exit if problem is solved
24:        end if
25:     end while
26:     # Step 2: Compute splitting metric
27:     ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z_{\text{new}}}\leftarrow{\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}h}({\color[rgb]{0.3671875,0.7890625,0.3828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.3671875,0.7890625,0.3828125}R_{\text{new}}})$
28:     ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z_{\text{last}}}\leftarrow{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}.\text{decomposition}[-1].{\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z}\ \textbf{if}\ {\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}.\text{decomposition}\neq\varnothing\ \textbf{else}\ -\infty$
29:     # Step 3: Split the step with the higher metric
30:     bsplit new step $\leftarrow$ ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z_{\text{new}}}\geq{\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z_{\text{last}}}$
[…]
32:     ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{1}},{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{2}}\leftarrow{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}f}({\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{\text{target step}}},{\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}\alpha})$
33:     if ${\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{1}}=\text{None}\ \textbf{or}\ {\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{2}}=\text{None}$ then
[…]
35:        Return ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}$vartriangleright Exit if we cannot do a finer split
36:     end if
37:     if $\textbf{b}_{\text{split new step}}$ then
[…]
39:        if ${\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z_{\text{new}}}<{\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}\theta}$ then
[…]
41:           Return ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}$vartriangleright Exit if all metrics are smaller than precision
42:        end if
43:     else
44:        ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}.\text{decomposition}[-1]\leftarrow\{\text{step\_str}:{\color[rgb]{0.265625,0.46484375,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.46484375,0.66796875}y_{1}},\text{metric}:{\color[rgb]{0.140625,0.6640625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.140625,0.6640625,0.8828125}z_{\text{last}}}\}$vartriangleright Split last step
45:     end if
46:  end while
47:  Return ${\color[rgb]{0.9921875,0.5546875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.9921875,0.5546875,0.234375}D}$
Appendix C Ablation studies
C.1 Ablation on Temperature
We conduct an ablation study to analyze the effects of temperature on DISC and BoN. Temperature controls the randomness of token sampling in autoregressive models, influencing both exploration and consistency. Higher temperatures encourage more diverse outputs, whereas lower temperatures yield more deterministic generations. To examine its impact, we evaluate DISC and BoN on a 100-problem subset of APPS (the first 100 problems) using gpt-4o-mini.
Figure 17 presents the Pass@token scaling curve for DISC across different temperatures. The results indicate that lower temperatures lead to improved performance, as DISC benefits from more deterministic step selection. Unlike BoN, which relies on broad solution sampling, DISC dynamically refines steps, making stable token probabilities advantageous.
Figure 18 illustrates the frequency of actual partitions made by DISC at different temperatures. As temperature increases, the number of partitions fluctuates more, suggesting that high temperature introduces instability in step selection. Lower temperatures provide more structured decomposition, reducing unnecessary subdivisions.
In Figure 19, we visualize the mean reward per step. The trend shows a linear increase in reward as step number grows, demonstrating that deeper decomposition results in progressively better solutions. This reinforces that DISC effectively allocates computation towards refining difficult steps.
The mean standard deviation per step is shown in Figure 20. Lower temperatures yield lower standard deviations, confirming that DISC benefits from reduced variability in sample quality. This consistency allows for more reliable prioritization of difficult steps, enhancing overall inference efficiency.
For comparison, Figure 22 and Figure 21 display Pass@token and Pass@k scaling curves for BoN across different temperatures. Unlike DISC, BoN achieves peak performance at a temperature around 0.6-0.8, balancing diversity and consistency. Higher temperatures increase exploration but degrade precision, while lower temperatures hinder sample diversity, reducing the probability of obtaining high-quality completions.
These findings highlight the fundamental difference between DISC and BoN: DISC benefits from lower variance and stable decomposition, while BoN relies on broader exploration facilitated by moderate temperature settings. As a result, optimal temperature settings differ significantly between these methods, with DISC favoring deterministic sampling and BoN requiring a balance between diversity and coherence.
Figure 17: Pass@token scaling curve for different temperatures on APPS using gpt-4o-mini. The lower the temperature, the stronger the DISC performance.
Figure 18: Partition frequency of DISC with different temperatures on APPS using gpt-4o-mini
Figure 19: Mean reward per step of DISC with different temperatures on APPS using gpt-4o-mini. The mean reward scales linearly with step number.
Figure 20: Mean standard deviation per step of DISC with different temperatures on APPS using gpt-4o-mini. Lower temperature means lower average standard deviation.
Figure 21: Pass@k scaling curve for different temperatures on APPS using gpt-4o-mini for BoN. A temperature around 0.6-0.8 leads to the best performance and balance between diversity and consistency.
Figure 22: Pass@token scaling curve for different temperatures on APPS using gpt-4o-mini for BoN. A temperature around 0.6-0.8 leads to the best performance and balance between diversity and consistency.
C.2 Ablation on Priority Metric $h$
We analyze the effect of different priority metrics on DISC performance. We evaluate DISC using the first 200 competition-level APPS problems with gpt-4o-mini, setting the temperature to 0.8 for all experiments. The priority metric determines which steps are refined during recursive decomposition, impacting both efficiency and final solution quality.
Figure 23 presents a token-level comparison of different priority metrics. Both DISC-Q and DISC-Z significantly outperform random selection and their inverse counterparts, demonstrating the importance of prioritizing high-value steps.
Figure 24 illustrates the partition frequency under different priority metrics. We observe that effective metrics such as DISC-Q and DISC-Z lead to fewer, more meaningful partitions, whereas suboptimal strategies result in excessive, redundant partitioning.
The relationship between mean reward and step number is shown in Figure 25. All tested metrics exhibit a strong correlation between increasing step depth and mean reward, indicating that decomposition progressively refines solutions. However, DISC-Q and DISC-Z achieve higher reward gains at earlier stages, suggesting that they prioritize the most impactful refinements.
Finally, Figure 26 reports the standard deviation of rewards per step. Lower standard deviation suggests more stable solution quality, a property that DISC-Q and DISC-Z maintain better than random selection methods. This highlights their effectiveness in identifying and refining challenging steps efficiently.
Overall, these results confirm that choosing an appropriate priority metric is crucial for DISC. While DISC-Q and DISC-Z consistently enhance inference efficiency and quality, random or inverse strategies lead to poorer performance due to misallocation of compute resources.
Figure 23: Token level comparison of different priority metrics on DISC in the APPS setting with gpt-4o-mini. Both Q and Z based priority metrics perform well.
Figure 24: Partition frequency of DISC with different priority metrics on APPS using gpt-4o-mini
Figure 25: Mean reward per step of DISC with different priority metrics on APPS using gpt-4o-mini. All metrics display strong correlation between step depth and the mean reward.
Figure 26: Mean standard deviation per step of DISC with different priority metrics on APPS using gpt-4o-mini. All metrics display correlation between step depth and the standard deviation.
C.3 Model Ablation
We investigate how different LLMs perform when used with DISC on 200 competition-level APPS problems, given a sample budget of 30. The groundtruth reward model was used to evaluate correctness, and all models were set to a temperature of 0.8. Due to the challenging nature of the benchmark, open-source models struggled to achieve strong performance independently. However, when paired with DISC, their performance significantly improved.
Figure 27 presents the Pass@token scaling curve for open-source models using DISC. The results demonstrate that DISC substantially enhances the capabilities of these models, closing the gap between them and proprietary alternatives.
Figure 28 visualizes the partition frequency of DISC with different open-source models. Compared to their standalone performance, the use of DISC led to more structured and effective decomposition, highlighting its adaptability to different architectures.
The mean reward per step is shown in Figure 29. Similar to prior findings, we observe that deeper decomposition leads to increasingly higher rewards. Notably, even lower-capacity models benefit from DISC ’s ability to iteratively refine their solutions.
Finally, Figure 30 presents the mean standard deviation per step. With DISC, the variance in performance is significantly reduced, resulting in more stable and reliable inference.
Overall, these findings emphasize that DISC is a robust framework capable of enhancing inference performance across diverse LLMs, particularly those with limited standalone capabilities.
Figure 27: Pass@token scaling curve for open source models with DISC on APPS. DISC also demonstrates strong performance gains with open source models.
Figure 28: Partition frequency of DISC with open source models on APPS
Figure 29: Mean reward per step of DISC with open source models on APPS
Figure 30: Mean standard deviation per step of DISC with open source models on APPS
C.4 Ablation on partition fraction $\alpha$
We include some more analysis on the partition fraction here.
Figure 31: Token level comparison of different DISC splitting fraction $\alpha$ on APPS competition level. $0.15\leq\alpha\leq 0.25$ seems to be optimal.
Appendix D Experimental results extended
D.1 APPS
Figure 32: Sampling frequency of each step averaged over the problems on APPS with gpt-4o-mini. DISC seems to have a preferences for spending more compute on earlier found steps.
D.2 Additional examples
Below is another computed decomposition for the same problem as the one shown in the main text:
Let $x$ be the length and $y$ be the width of the rectangle. Since the perimeter is 24 inches, we have $122x+2y=24\Rightarrow x+y=$ 12 . Therefore, we must maximize the area $xy$ subject to this constraint.
We use the method of Lagrange multipliers. The Lagrangian is
${\color[rgb]{0.265625,0.00390625,0.328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.00390625,0.328125}\mathcal{L}(x,y,\lambda)=xy-\lambda(x+y-12).}$
We differentiate with respect to $x,y,$ and $\lambda$ to obtain
${\color[rgb]{0.265625,0.00390625,0.328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.00390625,0.328125}\frac{\partial\mathcal{L}}{\partial x}=y-\lambda=0,\qquad\frac{\partial\mathcal{L}}{\partial y}=x-\lambda=0,\qquad\frac{\partial\mathcal{L}}{\partial\lambda}=x+y-12=0.}$
We find that $x=y=\lambda$, so $x=y=\frac{12}{2}=6.$ Therefore, the maximum area of the rectangle is $6\cdot 6=\boxed{36}.$
D.3 APPS with self-generated validation tests
We examine DISC performance on APPS when using self-generated validation tests. All methods utilized the same set of self-generated validation tests to ensure fair comparisons. Each problem received 5-10 validation tests, with the exact number determined dynamically by the LLM. We evaluated a subset of 100 APPS problems, generating samples until the sample budget was exhausted or a correct solution was found.
Figure 33 illustrates the Pass@token scaling curve, showing that DISC maintains strong scaling performance in this setting, though at a slightly lower rate compared to ground-truth verification.
Figure 34 and Figure 35 compare actual and planned partition frequencies, respectively. The results indicate that DISC continues to make structured decompositions even with self-generated validation, preserving its efficiency.
The mean reward per step, shown in Figure 36, follows a similar trend as in previous experiments, reinforcing that DISC effectively allocates compute resources for iterative refinement.
Lastly, Figure 37 demonstrates that DISC maintains lower standard deviations in performance, indicating stable quality improvements across steps.
Figure 33: Token level comparison of different decomposition methods on APPS with gpt-4o-mini and self-generated validation tests. DISC still scales better than other methods in this setting, albeit at a lower rate.
Figure 34: Actual partition frequency of different decomposition methods on APPS with gpt-4o-mini and self-generated validation tests.
Figure 35: Planned partition frequency of different decomposition methods on APPS with gpt-4o-mini and self-generated validation tests.
Figure 36: Mean reward per step of different decomposition methods on APPS with gpt-4o-mini and self-generated validation tests.
Figure 37: Mean standard deviation different decomposition methods on APPS with gpt-4o-mini and self-generated validation tests.
D.4 MATH500
Completions for MATH500 include both the reasoning steps and the final answer. Since MATH500 contains more problems than APPS200 and MATH problems tend to be relatively easier, solution quality saturates quickly. Therefore, we use a lower sample budget of 10 for these experiments.
Figure 38 presents the Pass@k performance for different decomposition methods on MATH500. We observe that all decomposition-based approaches achieve similar Pass@k performance, consistently outperforming BoN. This indicates that the structured nature of MATH problems allows multiple decomposition strategies to be effective.
Despite similar Pass@k results, the true advantage of DISC lies in its token efficiency, as shown in Figure 6. DISC significantly reduces the number of tokens required to reach correct solutions compared to alternative methods, demonstrating its ability to allocate computational effort efficiently in mathematical reasoning tasks.
Additionally, we analyze the partitioning behavior of DISC on MATH500. Figure 39 illustrates the actual partition frequency for different decomposition methods. The planned partitioning behavior, shown in Figure 40, further highlights how DISC effectively balances exploration and refinement.
Finally, we present the mean standard deviation per step in Figure 41. Lower variance suggests that DISC produces more stable and reliable decompositions over multiple runs, reinforcing its robustness in both mathematical and program synthesis domains.
Figure 38: Pass@k performance comparison for different decomposition methods on MATH500. DISC consistently outperforms BoN across different sampling budgets.
Figure 39: Observed partition frequency of different decomposition methods on MATH500. DISC effectively segments problems into meaningful subcomponents.
Figure 40: Planned partitioning strategy of different decomposition methods on MATH500. DISC ’s structured approach leads to more efficient problem breakdowns.
Figure 41: Mean standard deviation per step for different decomposition methods on MATH500. Lower variance in DISC suggests more stable and reliable problem-solving steps.
D.5 LiveCodeBench
We evaluate DISC on LiveCodeBench, a benchmark designed for code generation tasks with a focus on real-world software development challenges. LiveCodeBench presents a unique set of problems requiring both reasoning and structured decomposition, making it a suitable testbed for evaluating DISC’s ability to refine and improve intermediate steps.
Figure 42 shows the Pass@k comparison of different decomposition methods on LiveCodeBench. DISC consistently scales better than other decomposition methods, highlighting its ability to refine intermediate steps more effectively in complex coding scenarios.
Figure 43 illustrates the observed partition frequency of different decomposition methods. The structured approach of DISC results in well-balanced decomposition across steps, reducing unnecessary partitioning while maintaining sufficient granularity for improved solution refinement.
Figure 44 displays the planned partition frequency across methods. DISC dynamically determines the most effective partitions based on the evolving problem state, leading to more targeted and efficient decompositions.
Finally, Figure 45 presents the mean standard deviation per step across decomposition methods. Lower variance in DISC suggests that it produces more stable and reliable decompositions, reinforcing its robustness for solving LiveCodeBench problems.
Figure 42: Pass@k performance comparison for different decomposition methods on LiveCodeBench. DISC consistently outperforms other methods in structured problem refinement.
Figure 43: Observed partition frequency of different decomposition methods on LiveCodeBench. DISC effectively balances problem segmentation while avoiding excessive partitioning.
Figure 44: Planned partitioning strategy of different decomposition methods on LiveCodeBench. DISC dynamically adapts its partitioning to optimize search efficiency.
Figure 45: Mean standard deviation per step for different decomposition methods on LiveCodeBench. Lower variance in DISC suggests more stable and reliable problem-solving steps.
Appendix E Search and scaling
E.1 Monte Carlo Tree Search (MCTS)
Monte Carlo Tree Search (MCTS) is a widely used algorithm for sequential decision-making in large search spaces, particularly in applications such as game playing, planning, and inference scaling. The algorithm builds a search tree incrementally by simulating different sequences of actions and updating estimates of state quality. A key advantage of MCTS is its ability to balance exploration (discovering new states) and exploitation (refining promising ones) using a data-driven search process. The MCTS pipeline consists of four fundamental steps: selection, expansion, simulation, and backpropagation.
E.1.1 Selection
Starting from the root node representing the current state $\boldsymbol{s}$, MCTS iteratively traverses the search tree by selecting child nodes based on a selection policy. The most commonly used selection criterion is the Upper Confidence Bound for Trees (UCT), which balances exploration and exploitation:
$UCT(\boldsymbol{s},\boldsymbol{d})=\hat{Q}(\boldsymbol{s},\boldsymbol{d})+c\sqrt{\frac{\ln\left(\sum_{\boldsymbol{b}}n(\boldsymbol{s},\boldsymbol{b})\right)}{n(\boldsymbol{s},\boldsymbol{d})}},$ (1)
where $\hat{Q}(\boldsymbol{s},\boldsymbol{d})$ represents the estimated value of selecting action $\boldsymbol{d}$ from state $\boldsymbol{s}$, $n(\boldsymbol{s},\boldsymbol{d})$ is the visit count for this action, and $c$ is a hyperparameter controlling the trade-off between exploring new actions and favoring those with high past rewards.
E.1.2 Expansion
Once a leaf node (a previously unexplored state) is reached, the algorithm expands the tree by adding one or more new nodes. These new nodes represent potential future states $\boldsymbol{s}^{\prime}$ generated by sampling an action $\boldsymbol{d}$ from a predefined policy. This step broadens the search space and allows MCTS to evaluate new possibilities.
E.1.3 Simulation
Following expansion, the algorithm conducts a simulation (or rollout) from the newly added state. This step involves generating a sequence of actions according to a predefined policy until reaching a terminal state or an evaluation horizon. The outcome of the simulation, denoted as $v(\boldsymbol{s}^{\prime})$, provides an estimate of the quality of the new state. Depending on the application, this could represent a game result, an optimization score, or an inference accuracy metric.
E.1.4 Backpropagation
The final step involves propagating the results of the simulation back up the search tree to refine the estimated values of prior states and actions. Each node along the trajectory $\tau=[\boldsymbol{s}_{0},\boldsymbol{d}_{1},\boldsymbol{s}_{2},\dots,\boldsymbol{s}_{-1}]$ is updated iteratively:
$\hat{Q}(\boldsymbol{s}_{i},\boldsymbol{d}_{i+1})^{(t+1)}\leftarrow(1-\alpha_{n})\hat{Q}(\boldsymbol{s}_{i},\boldsymbol{d}_{i+1})^{(t)}+\alpha_{n}\max\{\hat{Q}(\boldsymbol{s}_{i},\boldsymbol{d}_{i+1})^{(t)},\hat{Q}(\boldsymbol{s}_{i+1},\boldsymbol{d}_{i+2})^{(t+1)}\},$ (2)
where $\alpha_{n}$ is a learning rate that depends on the visit count, and the maximum function ensures that the best-performing trajectories are emphasized.
MCTS has been widely adopted in inference scaling techniques due to its ability to efficiently allocate computational resources, focusing more on high-reward states while avoiding unnecessary exploration of unpromising regions. In later sections, we explore how MCTS can be combined with dynamic decomposition to further optimize inference scaling.
E.1.5 Combining Dynamic Decomposition with MCTS
MCTS can be enhanced by integrating dynamic decomposition, where each node in the search tree represents a decomposition of the problem into steps. Instead of treating states as atomic decisions, we recursively decompose reasoning steps, dynamically adjusting granularity based on difficulty.
In this framework:
  • Each node in the MCTS tree represents a partial decomposition of the problem, with child nodes corresponding to alternative step partitions.
  • Branching occurs by generating candidate next steps using dynamic decomposition, allowing finer steps for complex regions while maintaining efficiency for simpler ones.
  • The selection step prioritizes nodes that represent more promising decompositions, dynamically refining challenging areas through recursive subdivision.
  • The backpropagation step ensures that decompositions leading to high-quality solutions are reinforced, helping the search tree converge toward optimal inference paths.
By integrating dynamic decomposition with MCTS, we efficiently allocate compute to the most critical reasoning steps, improving inference quality while maintaining computational efficiency.
E.2 Beam Search
Beam search is a heuristic search algorithm commonly used in inference tasks where computational efficiency is a priority. Unlike exhaustive search methods, beam search maintains only the top $k$ best candidates at each step, making it an effective strategy for structured prediction problems and sequential decision-making.
At each iteration:
  • The algorithm selects the $k$ most promising partitions from the previous step based on an evaluation metric.
  • Each selected partition is expanded by generating possible next-step samples.
  • The newly generated partitions are ranked, and only the top $k$ candidates are retained for the next iteration.
  • This process continues until a stopping criterion is met, such as reaching a predefined depth or finding a sufficiently high-quality solution.
Beam search provides a computationally efficient way to explore structured solution spaces while maintaining high-quality search trajectories. By integrating beam search with dynamic decomposition, we ensure that inference computation is allocated efficiently, focusing on the most promising reasoning paths at each step.
E.3 Additional Results and Analysis
Experiments comparing different search methods were conducted on a 100-problem subset of the APPS dataset (first 100 problems) using GPT-4o-mini. All methods used a temperature of 0.2, with $\alpha=0.15$, Q priority metric, and $\sigma=1.0$.
Token-level comparison: As shown in Figure 46, MCTS scales best among the tested methods, demonstrating superior efficiency in identifying promising partitions. Greedy search follows closely, while beam search exhibits the slowest scaling.
Partition frequency analysis: Figure 47 reveals that greedy search explores to greater depths within the same sampling budget. This suggests that greedy search prioritizes deep refinements, whereas MCTS and beam search balance depth with breadth.
Step variance analysis: Figure 48 illustrates that all search methods display decreasing standard deviation with increasing search depth. This trend indicates that deeper searches converge towards stable, high-quality partitions, reinforcing the benefits of dynamic decomposition.
These results highlight the trade-offs between search methods: MCTS offers robust exploration-exploitation balance, greedy search favors depth-first refinement, and beam search provides a structured yet computationally constrained approach. The integration of dynamic decomposition further enhances these search strategies by adaptively allocating computational resources to critical reasoning steps.
Figure 46: Token level comparison of different decomposition search methods combined with DISC on APPS with gpt-4o-mini. MCTS scales best, followed by greedy search, followed by beam search.
Figure 47: Actual partition frequency of different decomposition search methods combined with DISC on APPS with gpt-4o-mini. Greedy is able to search to higher depths given the same sampling budget.
Figure 48: Mean standard deviation of different decomposition search methods combined with DISC on APPS with gpt-4o-mini. All search methods display decreasing standard deviation with search depth.